OpenStudy (anonymous):

(08.03)A set of equations is given below: Equation S: y = x + 9 Equation T: y = 2x + 1 Which of the following steps can be used to find the solution to the set of equations? x = 2x + 1 x + 9 = 2x x + 1 = 2x + 9 x + 9 = 2x + 1

OpenStudy (texaschic101):

any thoughts ?

OpenStudy (texaschic101):

try substitution and what do u get ?

OpenStudy (anonymous):

For example, consider the equations: x + y = 8 and x – y = 2. x = 1, y = 7 is a solution of the first equation but not the second. x = 4, y = 2 is a solution of the second equation but not the first. x = 5, y = 3 is a solution of both equations at the same time, so it is the solution of these simultaneous equations.

OpenStudy (anonymous):

'You might do this problem in your head. Sean’s daughter is 7, so “four times his daughter’s age” is 28, and 5 years added to that is 33. Sean is 33. If you solved the problem like that, you used a simple substitution—you substituted in the value “7” for “his daughter’s age.” You learned in the second part of the problem that “his daughter is 7.” So substituting in a value of “7” for “his daughter’s age” in the first part of the problem was okay, because you knew these two quantities were equal. Let’s look at a simple system of equations that can be solved using substitution.

OpenStudy (error1603):

What?

OpenStudy (texaschic101):

y = x + 9....so sub in x + 9 in for y in the other equation

OpenStudy (anonymous):

thank you guys so much i got it now;0

OpenStudy (error1603):

Report them please guys

OpenStudy (anonymous):

why they are just trying to help i think and hope lol

OpenStudy (error1603):

Look at all that! That's not helping.

OpenStudy (anonymous):

true:0 why dont you report them